# What Do the Circuits Mean? A Knowledge Edit View

### v2

@@ 6.3 The complementary circuit is more susceptible to the model editing
After applying the circuit-aware knowledge editing, we query the negative log probabilities of the HierarchyData entities again, and the outcomes are presented in Figure 2.
Figure 2: Post-edit negative log probability of new target distribution of 50% circuits.
@@
Our metric for assessing edit performance is the negative log probability of the new target, where lower values indicate better editing. The figure displays results for circuits extracted using the ‘determiner noun agreement 1’ (dna1) dataset of BLiMP, circuits extracted using hierarchy data (h) along with their complementary circuits, a randomly extracted circuit alongside its complementary circuit, and the full model circuit. Further details about other datasets are provided in Appendix A. Notably, the complementary 50% circuit consistently outperforms the 50% circuit in terms of editing for both the dna1 and h circuits. Furthermore, the performance of the random circuit is similar to that of its complementary counterpart. These indicate that the complementary circuits, which are less task-relevant, are more susceptible to knowledge edits. Yet, the full model demonstrates the best editing performance, indicating that the circuit might still synergize with the complementary networks upon the injection of new knowledge.
@@
Similar to the pre-edit observations, the post-edit discrepancies between the circuits and their complements can be generalized to a linguistic task (dna1). This suggests that task-specific circuits might indicate the knowledge storage locations. In other words, altering the knowledge from these knowledge-rich circuits is more difficult than altering the knowledge from their complements. Interestingly, humans have similar "confirmation bias" behavior as well: it’s more difficult to change someone’s opinion once it’s formed.
@@
Our findings here can also provide a novel perspective about neural network plasticity (Lyle et al., 2023), but a further in-depth inquiry into the nature of circuit-specific plasticity is left to future works.
7 How large are the ideal "theoretical circuits"?
@@
In our previous experiments, we conducted experiments on circuits that comprised 50% of the total model parameters. To investigate potential disparities, we reduced our analysis to include circuits that accounted for 5% of the parameters. Figure 3 (a) illustrates the outcomes of editing the 5% vs. the 50% circuits, along with their corresponding complementary circuits, using HierarchyData. We continue to employ the negative log probability of the new targets to quantify the edit performance. As comparisons, Figure 3 (b) presents the negative log probability scores of the true targets before model edits. Detailed figures of 5% circuits derived from other datasets are presented in Appendix B.
@@
Figure 3 (a) shows that both the 5% and the 50% circuits are more hard to edit than their corresponding complementary circuits. Notably, the 5% circuit displays a higher level of such "confirmation bias" than the 50% circuit. This suggests that the 5% circuit, rather than the 50% circuit, may be closer to an idealized "theoretical circuit" where essential knowledge is concentrated.
@@
Now, if a theoretical circuit exists, would its size be smaller than 5% or between 5% and 50%?
@@
As Figure 3 (b) shows, the 5% circuit shows the best capturing of knowledge, followed by the 50% circuit, because the 50% circuit may contain non-knowledgeable "noise" parameters that mask out the circuit’s ability to demonstrate knowledge. Also shown in this figure, the complementary of 5% circuit is more knowledgeable than the complementary of 50% circuit, indicating that the 5% complementary circuit likely possesses certain key parameters directly associated with knowledge retention, whereas the 50% complementary circuit may include a surplus of less critical parameters. In other words, a "theoretical circuit", if it exists, would incorporate more than 5% but less than 50% of parameters.
@@
Table 2 provides additional evidence to how large the idealized circuits are. We replicate the experiments, this time extracting 15%, 25%, 35% parameters circuits while using the same hyperparameters. The table below shows the median post-edit negative log probabilities of the new target, combined with the results from the 50% and 5% circuits (higher values indicate worse performance). As expected, the edit performance of the extracted circuits declines as the circuit size decreases. If an idealized theoretical circuit exists that encapsulates all relevant knowledge in the model, it would likely have the worst edit performance due to strong ‘confirmation bias’. Therefore, we hypothesize that the 5% circuit may be closer to this idealized “theoretical circuit” than the 50% circuit. However, the exact size of this idealized circuit remains to be further investigated.
@@
Complementary circuit 0.0126 0.00999 0.00997 0.00985 0.00981
Table 2: Median post-Edit negative log probabilities of the new target across different percentage parameter circuits
Recent papers on circuit detection look for highly sparse circuits that usually contain less than 1% of the total parameters (Bayazit et al., 2023) or less than 3% of total edges (Hanna et al., 2024), depending on the tasks. Our observations indicate that a larger (i.e., >5%) subset of the networks on linguistic and knowledge-related tasks may also be worth of analysis.
Figure 3: Probability distribution of 5% circuit and 50% circuit. a. Post-edit probability of new target (top); b. Pre-edit probability of true target (bottom)
@@ 8 Do these circuits overlap?
Figure 4 features a heatmap that depicts the mask similarity across different tasks. Each entry in the heatmap represents the degree of overlap between 5% circuits from two datasets, measured relative to the base dataset listed in the row. For example, a value of ’0.137’ at the intersection of the second row and the first column suggests that the circuit derived from the ’dna1’ dataset overlaps with the ’h’ dataset circuit by 13.7% of the ’dna1’ circuit.
To create a baseline, we randomly select a 5% parameter circuit mask and calculate the overlap percentage of all parameters within the circuit. The results are presented in the table 3. The percentage is consistently around 5%.
dna1 dna2 h dnai1 dnawa1 dnawai1 e ra hhh
0.05026 0.05024 0.05013 0.05007 0.05021 0.05023 0.05001 0.04997 0.05007
Table 3: Percentage of parameter overlap between the circuits and a randomly selected circuit
The degrees of overlap between the circuits vary moderately across different dataset comparisons—indicating neither negligible nor overwhelming similarity. Interestingly, datasets that appear closely related, such as ’dna1’ and ’dna2’, sometimes display significant differences when viewed through the lens of the model’s mechanism. Notably, the ’ra’ (risk-averse) circuit shows greater overlaps with other circuits, suggesting that the risk-averse dataset may include terms that are broadly relevant across multiple contexts. Conversely, the ’e’ (extraversion) circuit demonstrates fewer overlaps, implying that the extraversion dataset likely features specialized terms that do not commonly occur in other datasets.
Additionally, we record the mask for the 50% circuit of 3 datasets and calculate the overlap percentage of all parameters within the circuit. Results are shown in Table 4. They each have approximately 60% overlap with one another, which is much higher than the overlap between the 5% circuits.
@@
h circuit 1 0.605 0.581
dna1 circuit 0.605 1 0.588
ra circuit 0.581 0.588 1
@@ 9 What Transformer modules are in the circuits?
This distribution of parameters is intriguing because it suggests that, while LayerNorm plays a crucial structural role, the computational responsibilities might be more evenly distributed or perhaps differently focused in attention and MLP layers. Notably, some studies found that the MLPs are critical repositories of knowledge within neural networks (Geva et al., 2020, 2022), and the attentions are critical for the mechanisms related to factual recall (Geva et al., 2023). Recently, Katz and Belinkov (2023) found that LayerNorm is related to the introduction of new tokens. There are also some recent debates about the exact localization of the knowledge (Hase et al., 2024; Niu et al., 2024). In contrast to prior works, our observations show that LayerNorm plays a critical role in the circuits.
Furthermore, the significant allocation of parameters to LayerNorm in the circuits raises questions about their specific function and impact on the overall network efficacy. Future analysis of LayerNorm could further advance our understanding.
@@ 10 Discussion
Limitations and recommendations for future researchers
We propose several recommendations for future researchers aiming to enhance model safety: 1) Deeper Analysis of Complementary Circuits: Further investigation into the role and characteristics of complementary circuits can provide insights into how they interact with primary circuits and affect overall model behavior. 2) Searching for Theoretical Circuits: Investigating the possibility of theoretical circuits that can perform well for every specific tasks. 3) Utilizing Other Editing Techniques: Exploring other techniques like ROME (Meng et al., 2022a) to edit complementary circuits without impacting primary circuits may help mitigate biases and improve model fairness. 4) Expanding Behavior Datasets: Creating more comprehensive datasets that capture a wider range of behaviors can enhance our understanding of the relationships between behaviors and the knowledge stored in LLMs. By addressing these recommendations, future research can contribute to creating safer and more reliable language models.
@@ 11 Conclusion
We propose an approach to study the functions of the automatically-extracted circuits as a holistic object, through the lens of knowledge editing. By examining the effects of circuit size, the overlap between circuits derived from different datasets, and the distinct roles of various layers within these circuits, we present a comprehensive analysis that enhances our understanding of LLM behavior.
We observe a "confirmation bias" behavior at the circuit level: the primary circuits extracted from classification tasks, which are presumed to contain related knowledge, are challenging to edit. In contrast, complementary circuits, which are considered less integral to knowledge storage, adapt more readily to new information. This suggests that editing the primary circuits involves significant challenges, akin to altering well-established beliefs.
The impact of the circuit size was also explored. We compare the results of 5% circuits and 50% circuits, and reveal that larger circuits may introduce noise, diluting their effectiveness in knowledge interpretation and storage. This finding suggests that maybe essential knowledge is densely concentrated in an ideal "theoretical circuit", which is likely close to the 5% circuit.
Additionally, we investigate into the overlap between circuits from different linguistic task datasets, neural network behavior datasets and knowledge datasets. We reveals moderate variations in overlap degrees, suggesting a balance between distinctiveness and commonality among circuits.
Finally, the roles of different types of layers within circuits are examined, revealing that layer normalization layers command a large share of parameters, particularly in smaller circuits. Attention and multi-layer perceptron (MLP) layers hold smaller but critical shares. This distribution challenges traditional views on knowledge storage in neural networks and suggests that MLP layers, in particular, may be vital knowledge repositories. The prominence of layer normalization layers underscores their potential impact on network stability and performance, highlighting the need for deeper investigation into how these layers influence the overall efficacy of LLMs.
In conclusion, our research provides a detailed understanding of how automatically-extracted circuits within LLMs perform in knowledge editing tasks. These findings support future explorations into model behaviors, bias removal, and interpretability. By addressing the open problems and recommendations outlined in our discussion, future research can contribute to creating safer, fairer, and more reliable language models.
@@
Figures 5 and 6 display the post-edit probability of the new target and the pre-edit probability of the true target for the BLiMP dataset, respectively. Similarly, Figures 7 and 8 show these distributions for the Model-Written Evaluations dataset.
@@
Figures 9 and 10 show the post-edit probability of the new target and the pre-edit probability of the true target for the BLiMP dataset, respectively. Figures 11 and 12 display the post-edit probability of the new target and the pre-edit probability of the true target for the Model-Written Evaluations dataset, respectively.

### v3

@@ 6.3 The complementary circuit is more susceptible to the model editing
After applying the circuit-aware knowledge editing, we query the negative log probabilities of the HierarchyData entities again, and the outcomes are presented in Figure 1 (right).
@@
Our metric for assessing edit performance is the negative log probability of the new target, where lower values indicate better editing. The figure displays results for circuits extracted using the “determiner noun agreement 1” (dna1) dataset of BLiMP, circuits extracted using hierarchy data (h) along with their complementary circuits, a randomly extracted circuit alongside its complementary circuit, and the full model circuit. Further details about other datasets are provided in Appendix A. Notably, the complementary 50% circuit consistently outperforms the 50% circuit in terms of editing for both the dna1 and h circuits. Furthermore, the performance of the random circuit is similar to that of its complementary counterpart. These indicate that the complementary circuits, which are less task-relevant, are more susceptible to knowledge edits. Yet, the full model demonstrates the best editing performance, indicating that the circuit might still synergize with the complementary networks upon the injection of new knowledge.
@@
Similar to the pre-edit observations, the post-edit discrepancies between the circuits and their complements can be generalized to a linguistic task (dna1). This suggests that task-specific circuits might indicate the knowledge storage locations. In other words, altering the knowledge from these knowledge-rich circuits is more difficult than altering the knowledge from their complements.
@@
Interestingly, humans have similar “confirmation bias” behavior, tending to reject hypotheses that do not agree with prior assumptions (Wason, 1960). Our findings here can also provide a novel perspective about neural network plasticity (Lyle et al., 2023), but a further in-depth inquiry into the nature of circuit-specific plasticity is left to future works.
Figure 2: Pre-edit (left) and post-edit (right) log-probability distribution of 5% circuit and 50% circuit
7 How large are “ideal circuits”?
Suppose that there exists a circuit that stores all knowledge needed to perform the task — let’s use the term “ideal circuit” to describe it — how large would it be?
In our previous experiments, we conducted experiments on circuits that comprised 50% of the total model parameters. To investigate potential disparities, we reduced our analysis to include circuits that accounted for 5% of the parameters, using HierarchyData.
@@ 7 How large are “ideal circuits”?
Figure 2 (left) presents the negative log probability scores of the true targets before model edits, and Figure 2 (right) illustrates the outcomes of editing the 5% vs. the 50% circuits, along with their corresponding complementary circuits. We continue to employ the negative log probability of the new targets to quantify the edit performance. Detailed figures of 5% circuits derived from other datasets are presented in Appendix B.
@@
Figure 2 shows that both the 5% and the 50% circuits are harder to edit than their corresponding complementary circuits. Notably, the 5% circuit displays a higher level of such “confirmation bias” than the 50% circuit. This suggests that the 5% circuit may be closer to an ideal circuit, than the 50% circuit.
@@
Now, if an ideal circuit exists, would its size be smaller than 5% or between 5% and 50%?
@@
As the pre-edit logprob distribution plot in Figure 2 (left) shows, the 5% circuit captures the knowledge the best, followed by the 50% circuit (perhaps because the 50% circuit may contain non-knowledgeable “noise” parameters that mask out the circuit’s ability to demonstrate knowledge). Also shown in this figure, the complementary of 5% circuit is more knowledgeable than the complementary of 50% circuit, indicating that the 5% complementary circuit likely possesses certain key parameters directly associated with knowledge retention, whereas the 50% complementary circuit may include a surplus of less critical parameters. In other words, an ideal circuit, if it exists, would incorporate more than 5% but less than 50% of parameters.
@@
Table 2 provides additional evidence toward the sizes of the idealized circuits. We replicate the experiments, this time extracting 15%, 25%, 35% parameters circuits while using the same hyperparameters. The table below shows the median post-edit negative log probabilities of the new target, combined with the results from the 50% and 5% circuits (higher values indicate worse performance). As expected, the edit performance of the extracted circuits declines as the circuit size decreases. If an ideal circuit exists that encapsulates all relevant knowledge in the model, it would likely have the worst edit performance due to strong “confirmation bias”. Therefore, we hypothesize that the 5% circuit may be closer to this “ideal circuit” than the 50% circuit. However, the exact size of this idealized circuit remains to be further investigated.
@@
Complementary circuit 1.26e-2 9.99e-2 9.97e-2 9.85e-2 9.81e-2
Table 2: Median post-edit negative log probabilities of the new target across circuits with different sizes
Recent papers on circuits detected highly sparse circuits that usually contain less than 1% of the total parameters (Bayazit et al., 2023) or less than 3% of total edges (Hanna et al., 2024) for replicating the model behavior. We hypothesize that while the circuits less than 1% explain the functions, the “ideal circuits” need more parameters that correspond to the knowledge used in the tasks. Also, considering that the knowledge localization might differ from the model editing (Hase et al., 2024), the roles of distinct parameters in the model would be an intriguing path to explore further.
@@ 8 Do these circuits overlap?
To create a baseline, we randomly select a 5% parameter circuit mask and calculate the overlap percentage of all parameters within the circuit. As shown in Table 3, the percentages constantly appear around 5% and only deviate by errors that are two magnitudes smaller. Mathematically, the expectation of a circuit with Xparameter overlaps with a random Xcircuit is X%.
Figure 3 features a heatmap that depicts the mask similarity across different tasks. Each entry in the heatmap represents the degree of overlap between 5% circuits from two datasets, measured relative to the base dataset listed in the row. For example, a value of 0.137 at the intersection of the second row and the first column suggests that the circuit derived from the dna1 dataset overlaps with the h dataset circuit by 13.7% of the dna1 circuit.
Task dna1 dna2 h dnai1 dnawa1 dnawai1 e ra hhh
Difference from 0.05 2.6e-4 2.4e-4 1.3e-4 0.7e-4 2.1e-4 2.3e-4 0.1e-4 -0.3e-4 0.7e-4
Table 3: Percentage of parameter overlap between each 5% circuit and a randomly selected circuit
The degrees of overlap between the circuits vary across different dataset comparisons — between 13% and 18% — which are at least twice the baseline of 5%. These indicate non-negligible similarities. Interestingly, datasets that appear closely related, such as dna1 and dna2, sometimes display significant differences when viewed through the lens of the model’s mechanism. Notably, the ra (risk-averse) circuit shows greater overlaps with other circuits, suggesting that the risk-averse dataset may include terms that are broadly relevant across multiple contexts. Conversely, the e (extraversion) circuit demonstrates fewer overlaps, implying that the extraversion dataset likely features specialized terms that do not commonly occur in other datasets.
Similarly, the 50% circuit of three datasets overlap with each other at levels significantly higher than the 50% expectation, as shown in Table 4.
@@
h circuit 1 0.61 0.58
dna1 circuit 0.61 1 0.59
ra circuit 0.58 0.59 1
@@ 9 What Transformer modules are in the circuits?
This distribution of parameters is intriguing because it suggests that, while LayerNorm plays a crucial structural role, the computational responsibilities might be more evenly distributed or perhaps differently focused in attention and MLP layers. Notably, some studies found that the MLPs are critical repositories of knowledge within neural networks (Geva et al., 2020, 2022), and the attentions are critical for the mechanisms related to factual recall (Geva et al., 2023). Recently, Katz and Belinkov (2023) found that LayerNorm is related to the introduction of new tokens. There are also some recent debates about the exact localization of the knowledge (Hase et al., 2024; Niu et al., 2024). Complementing prior works, our observations show that LayerNorm plays a critical role in the circuits related to the syntax and hierarchical knowledge entity tasks.
The significant allocation of parameters to LayerNorm in the circuits raises questions about their specific function and impact on the overall network efficacy. Future analysis of LayerNorm could further advance our understanding.
@@ 10 Discussion
Implications for model training
The understanding of the information structures in the circuits can lead to fine-tuning technologies. The neural network circuit is a particularly appropriate medium because the sparsity can lead to high efficiency, and the automatic circuit discovery algorithms allow the search of circuits that sacrifice model performance to the fewest possible extent. We plan to continue developing the training algorithms in future works.
Recommendations for safer models
We propose several recommendations for future researchers aiming to enhance model safety. First, further investigation into the role and characteristics of complementary circuits can provide insights into how they interact with primary circuits and affect overall model behavior. Second, investigating the possibility of ideal circuits that can perform well for every specific task. Third, exploring other techniques like ROME (Meng et al., 2022a) to edit complementary circuits without impacting primary circuits may help mitigate biases and improve model fairness. Fourth, creating more comprehensive datasets that capture a wider range of behaviors can enhance our understanding of the relationships between behaviors and the knowledge stored in LLMs. Along these paths, future research can contribute to creating safer and more reliable language models.
@@ 11 Conclusion
In this paper, we combine two model analysis techniques: knowledge edits and circuit analysis, to inspect the model’s knowledge structure. Our comprehensive experiments reveal novel properties in the automatically extracted circuits, including the resistance to knowledge edit, which we compare to the “confirmation bias”. Our experiments reveal distribution patterns of the knowledge within circuits, across the circuits, and the roles of the knowledge in safety behavior tasks. These findings support future explorations into model behaviors, bias removal, and interpretability. By addressing the open problems and recommendations outlined in our discussion, future research can contribute to creating safer, fairer, and more reliable language models.
12 Limitation
We used only the GPT-2 model. Its structure (Transformer decoder) is the backbone of the current LLMs, and the modules are widely used. We only investigated fine-tuning model edit, and alternative model editing methods can be explored. We investigated a handful of text classification problems including linguistic tasks and safety behavior tasks, but there can always be more.
@@ 12 Limitation
Figures 4 and 5 display the post-edit probability of the new target and the pre-edit probability of the true target for the BLiMP dataset, respectively. Similarly, Figures 6 and 7 show these distributions for the Model-Written Evaluations dataset.
@@
Figures 8 and 9 show the post-edit probability of the new target and the pre-edit probability of the true target for the BLiMP dataset, respectively. Figures 10 and 11 display the post-edit probability of the new target and the pre-edit probability of the true target for the Model-Written Evaluations dataset, respectively.